what expression can be used to find the measure of a tank capacity in cubic feet if it measure 2ft by 4 ft by 3 ft?
Answer by rapaljer(4671)

Volume of a rectangle = L*W*H

V=2*4*3 = 24 cubic feet.

For additional explanation on this topic, please see my own website.
The best way to find it is to use the easy-to-remember and easy-to-spell link www.mathinlivingcolor.com. At the bottom of the page there is a single link that takes you to my Homepage. Near the bottom of my Homepage, look for the link "Topics from Geometry." Choose the link "Perimeter, Area, and Volume." Here you will find my own non-traditional explanation that my own students used to find much easier to understand than most of the other textbooks being used. I also have lots of examples, exercises, and answers.

To contact me please use my Email address at rapaljer@seminolestate.edu.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus